cpuld someone check this for me.Thanks in advance for your help.
solve the equation by completing the square
x^2+8x-13=0 i got x=-4+sqrt 29,-4- sqrt 29

Answer by stanbon(75887)

solve the equation by completing the square
x^2+8x-13=0
x^2+8x + 16 = 13+16
(x+4)^2 = 29
x+4 = sqrt(29) or x+4 = -sqrt(29)
x = -4 +sqrt(29) or x = -4-sqrt(29)
===================================
Cheers,
Stan H.